OpenStudy (anonymous):

Suppose that you are on a planet, that has an average radius 1*10^9 and average density 7.4 kg/m^3. Assuming that it is a perfect sphere, find the mass of the planet.

OpenStudy (anonymous):

* 1* 10^9 meters

OpenStudy (espex):

Well you know that \[density = \frac{mass}{volume}\] and that the volume of a sphere is \[\frac{4}{3} \pi r^3\] so if you rearrange that you get \[\frac{density}{volume}=mass\] which should yield an equation of \[\frac{3*density}{4\pi r^3}=mass\]

OpenStudy (anonymous):

Where does the 3 * density come into play?

OpenStudy (espex):

You are dividing your density by 4/3 in the volume equation.

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

22.20 kg/ m^3 / 1.257 * 10^16 km = 1.766 * 10 kg/m^2 is what I came out with

OpenStudy (espex):

Your meters should have all canceled out since mass is just a measure of kg.

OpenStudy (anonymous):

So its just 1.766 * 10 kg